# The occurrence of small, short-period planets younger than 200 Myr with TESS

### 05 Mar 2024 v1

@@ 4.2 Box Least Squares Transit Detection
With the detrended light curves, we perform a Box Least Squares (BLS; Kovács et al., 2002) search to identify planetary transits in the TESS light curves. We sample 100,000 periods linearly spaced between 0.5 and 20 days and 500 transit durations linearly sampled between 0.008 times and 0.08 times of each tested orbital period.
4.3 Vetting Threshold Crossing Events
@@
We identify any signal with a calculated Signal-to-Pink Noise (Hartman & Bakos, 2016) 8 to be a threshold crossing event (TCE). For each individual TCE, we perform a visual inspection to vet for false positives resulting from both astrophysical events and detrending artefacts. Each candidate event is inspected to determine if the signal that triggered a TCE is transit-shaped. Further, we require a minimum of three transit events in order to determine a period.
@@
We search for the presence of secondary eclipses and differences in the depths between even and odd transits to determine if the transit is from an obvious eclipsing binary (EB). For identified TCEs, we fit for the transit depth of even and odd transits respectively to rule out EB scenarios. To measure the transit depth, we fix all parameters except for transit depth, . The transits are modeled using batman (Kreidberg, 2015). We explore the best-fit transit depth and posteriors with a Markov chain Monte Carlo run implemented with emcee (Foreman-Mackey et al., 2013). We require the difference in our calculated odd and even transit depths to be to not be classified as an EB, though no reported candidate has an odd-even difference in our sample. We visually vet for secondary eclipses by phase-folding the detrended TESS light curves at the derived TCE period and transit time. The entire phase folded light curve is visually inspected for signs of secondary eclipses. Targets exhibiting eclipsing binary-like secondary eclipses, including those in eccentric orbits, are excluded as planet candidates.
@@
We check for obvious nearby eclipsing binaries (NEBs) by comparing the pixel-by-pixel light curves centered around the target star to determine if the signal that triggered the TCE is consistent with being on target. We then perform a centroid analysis with the TESS FFIs to create difference images using the transit-diffImage codebase333https://github.com/stevepur/transit-diffImage following the techniques described in Bryson et al. (2013). For each candidate at each sector, we produce a difference image between the in-transit and out-of-transit pixel stamps. We calculate the centroid of the target star via a TESS Pixel Response Function (PRF) fit to the difference image. The calculated centroid is compared against the expected centroid of the target star based on a catalog projection. Some offset in the difference image is expected if nearby stars are present, as the centroid naturally shifts away from the dimming star during transit. We flag any centroid offset pixel for visual inspection to determine if the high centroid offset is due to an astrophysical false-positive.
@@
Remaining candidates were then checked for planetary multiplicity. We masked the initial signal that triggered the TCE in the non-detrended light curves and then repeated the detrending and transit search (see Section 4.1 and 4.2). Any signals that pass our signal-to-noise criteria are visually vetted in the same method as mentioned above. We additionally check for aliasing of the original signal recovered. All remaining candidates are compared to the list of TOIs and known planets.
@@
Our stellar sample contained 26 TOIs (see Table 2), 22 of which were within our search parameters (0.5 days 20 days and ), as well as 2 previously previously confirmed transiting planets from K2. Of these, 16 have both FFI light curves and 2-minute cadence light curves (see Table 2), as well as the two K2 planets. Of the 22 known TOIs within our survey parameters, 20 are either planet candidates (PC) or confirmed planets, the remaining 2 have been ruled as either false-alarms or false-positives by TESS Follow-up Observing Program Working Groups (SG1 and SG2).
@@
We recovered all known TOIs in our planet search using the SPOC light curves, including 15 planet candidates within our search parameters and the known planets V1298 Tau c and d. Our QLP planet search recovered all 20 TOIs within our parameters but was unable to recover V1298 Tau c and d. Our pipeline identified a signal corresponding to V1298 Tau c, with a signal-to-pink-noise of 6.29, which did not trigger a TCE. Upon visual inspection of the FFI light curves for V1298 Tau, the long cadence of the FFIs and the heightened stellar activity subsequently diluted the transit signal in our detection pipeline.
@@
We identified new planet candidates TIC 88785435.01, TIC 150070085.01, TIC 434398831.01, and TIC 434398831.02 (Vach et al. in prep) in our QLP search. Out of the new planet candidates, only TIC 150070085.01 is also observed by SPOC data and is independently identified by our SPOC search. These candidates are described in detail in Appendix C.
| TIC | Name | R [] | [days] | Pipeline | [K] | Age [Myr] | Association | False Positive Rate |
@@
| 299798795 | TOI-1224.01 | 1.900.050 | 4.18 | SPOC & QLP | 3390 | 178 | MELANGE-5 | 0 |
| 299798795 | TOI-1224.02 | 3.000.14 | 17.95 | SPOC & QLP | 3390 | 178 | MELANGE-5 | 0 |
@@
| 384984325 | TOI-6109.01 | 4.510.46 | 8.54 | SPOC & QLP | 5760 | 55 | Theia133 | 0 |
| 384984325 | TOI-6109.02 | 2.780.32 | 5.70 | SPOC & QLP | 5760 | 55 | Theia133 | 0 |
| 410214986 | DS Tuc Ab | 5.700.17 | 8.14 | SPOC & QLP | 5410 | 40 | Tuc-Hor | 0 |
| 441420236 | AU Mic b | 4.200.20 | 8.46 | SPOC & QLP | 3590 | 25 | Pic | 0 |
| 441420236 | AU Mic c | 3.240.16 | 18.86 | SPOC & QLP | 3590 | 25 | Pic | 0 |
| 460205581 | TOI-837 b | 8.601.0 | 8.32 | SPOC & QLP | 6100 | 35 | IC2602 | 0 |
| 464646604 | HIP 94235 b | 3.000.21 | 7.71 | SPOC & QLP | 5990 | 120 | AB Dor | 0 |
@@
| 257605131 | TOI-451 b | 1.86 | SPOC & QLP | 5480 | 120 | Pisces-Eridanus | 0 | |
| 257605131 | TOI-451 c | 9.19 | SPOC & QLP | 5480 | 120 | Pisces-Eridanus | 0 | |
| 257605131 | TOI-451 d | 16.36 | SPOC & QLP | 5480 | 120 | Pisces-Eridanus | 0 | |
@@
| 46631742 | TOI-5358.01 | 2.610.14 | 2.66 | SPOC & QLP | 4640 | 134 | Theia369 | 0.011260.00088 |
| 157081737 | TOI-6095.01 | 1.870.16 | 2.16 | SPOC & QLP | 5260 | 106 | Theia301 | 0.12040.0020 |
@@
| 64837857 | TOI-6550.01 | 14.1 | 20.3 | SPOC & QLP | 6740 | 43 | Theia 117 | |
| 86951294 | TOI-4596.01 | 2.72 | 4.12 | SPOC & QLP | 5600 | 55 | Theia 133 | |
| 166527623 | HIP 67522 b | 10.1 | 6.96 | SPOC & QLP | 5770 | 17 | Sco-Cen | |
@@
| 117689799 | TOI-3504.01 | 14.2 | 2.16 | SPOC & QLP | 5860 | 99 | Theia 254 | |
| 133505138 | TOI-4359.01 | 2.47 | 2.436 | SPOC & QLP | 6530 | 91 | Theia 215 | |
| 164150539 | TOI-1372.01 | 8.22 | 6.16 | SPOC & QLP | 5770 | 100 | Theia 312 | |
| 294500964 | TOI-2550.01 | 10.5 | 9.07 | SPOC & QLP | 5740 | 83 | Theia 211 | |
| 457939414 | TOI-1990.01 | 11.8 | 3.78 | SPOC & QLP | 6080 | 35 | Theia 92 | |
David et al. (2019); Thia et al. in prep; Wood et al. (2023); in prep; Newton et al. (2019); Plavchan et al. (2020); Bouma et al. (2020); Zhou et al. (2022); This work; Vach et al. in prep; Newton et al. (2021); () Rizzuto et al. (2020); () Zakhozhay et al. (2022);
@@
4.4 False Positive Rate
Our planet detection pipeline recovered 14 confirmed planets within our SPOC parent sample, and 16 confirmed planets in the QLP parent sample. For planets with Confirmed Planet dispositions by the TESS Follow-up Observing Program Working Group, we adopt a false positive rate of zero, as these planets have undergone extensive spectroscopic and photometric follow-up observations to rule out false positive scenarios.
Five remaining planet candidates have not received sufficient follow-up to be categorized as verified or validated planets. As such, there exists a possibility that they are the result of astrophysical false positive scenarios, instrumental systematics, stellar variability, detrending artefacts, etc. There are not sufficient planet candidates around young stars from which we can derive a secure false positive rate estimate.
To attain false-positive estimates, we make use of the publically available TRICERATOPS (Giacalone et al., 2021) code to estimate the probability a transit signal is planetary in nature or the result of an astrophysical false positive. We explore the impact on the false positive rates when derived using blend analyses codes, including TRICERATOPS and MOLUSC (Wood et al., 2021). The false-positive rates from TRICERATOPS are adopted in our occurrence rate calculations hereafter and are typically 1-5% for a given candidate. When we adopt TRICERATOPS and MOLUSC together, the false-positive rate approaches 0% for all candidates, and its impact on the calculated occurrence rates is the same as if we assume all candidates are true planets. This upper-bound case is further explored in Appendix B. We do not attempt to differentiate the false positive rates between multi- and single-planet systems. We present our calculated false-positive rates in Table 2.
We adopt a false positive rate of 0% for TIC 434398831.01 and .02, which are presented in this work. These are characterized as confirmed planets as space- and ground-based follow-up have ruled out any false-positive scenarios (see Appendix C).
@@ 4.4 False Positive Rate
4.5 Candidates identified but not included in occurrence rate calculations
We identify 9 planets or planet candidates that are not included in our occurrence rate calculations. These are presented in Table 2. HIP 67522 b (, Rizzuto et al., 2020), HD 114082 (single transit, , Zakhozhay et al., 2022), and TOI-6550.01 (, days) were recovered in both our SPOC and QLP pipelines, but are not included in our occurrence rate calculations as they fall outside the bounds our survey investigates. We note that TOI-2519.01 (QLP only) and TOI-4596.01 (SPOC and QLP), which we recover in our survey, both show no Li features in ground-based follow-up observations. Upon visual inspection, their respective light curves are also not consistent with youth. We therefore exclude them from our occurrence rate calculations as they are likely not members of their respective clusters. TOI-4596 has an additional single transit event, but we are unable to constrain the period.
TOI-6555.01 is excluded as our vetting processes did not find it to be consistent with being on target. Additionally, TOI-6555.01 was not alerted until sector 65, which was not made available until after February 2023.
We identify two new additional candidates, TIC 36332984.01 and TIC 238395674.01, which both pass all of our initial vetting procedures, but we are unable to uniquely determine the orbital periods. TIC 238395674 is likely a multi-planet system, where we observe two transits of TIC 238395674.01 and an additional single transit of TIC 238395674.02. Therefore we do not include these candidates in our occurrence rate calculations.
Additionally, TOI-6715, previously identified as PATHOS 31 (Nardiello et al., 2020), has both QLP and SPOC data. However, as of February 2023 only the QLP data was available, therefore we only include TOI-6715 in our QLP occurrence rate calculations, but not in our SPOC occurrence rate calculations.
4.6 False positive TOIs not included in occurrence rate
@@
Follow-up observations by the TESS Follow-up Observing Program Working Groups (TFOP WG) rule five as nearby planet candidates (NPC), meaning the signal triggering the TCE is from a planet transiting a nearby star, or ambiguous planet candidates (APC). TFOP WG SG2 identified the remaining two as a nearby eclipsing binary (NEB) and an eclipsing binary (EB) respectively.
@@ 5 Injection Recovery Test
We perform a set of transit injection and recovery tests to test the completeness of the transit detection pipeline. We perform 1000 iterations of transit injections into each of the non-detrended 7154 QLP and 1927 SPOC light curves. The injected transits are modeled via batman (Kreidberg, 2015), with periods, , and radii, , sampled from log-uniform distributions ranging from 0.5 days to 20 days, and 1 to , impact parameters, , selected from a uniform distribution, between 0 and 0.8, and a transit time, , uniformly selected from a random phase. The models are resampled to 2-, 10-, and 30-minute cadences, and multiplied into the non-detrended QLP and/or SPOC light curves for each target. The injected light curve is then run through our detection pipeline as described in Section 4.
An injected planet is recovered if the BLS signal-to-pink-noise threshold is met and if the output ephemeris matches that of the injection. As with our planet detection pipeline, we require a TCE to have a signal-to-pink-noise . The detected TCE is also required to have its detected transit time and period from the BLS to match the injected time and period with a significance of and (see Coughlin et al. (2014) for definitions of and ).
We present our derived injection and recovery maps in Figure 6. These maps present our full series of injection and recovery tests summed across the entire parent stellar population of QLP and SPOC respectively. Our recoverability of larger planets (5 R) is consistent across the two surveys, but we see a sharper decline in recoverability in the region between 2-4 R for QLP compared to SPOC. This is consistent with our expectations due to the difference in stellar populations between the two samples. The on-average larger stellar radii in the QLP sample lead to more difficulties in recovering smaller planets.
We also investigate our susceptibility to false alarms with an inverted light curve test, as described in Appendix A, finding our manual vetting to be sufficient in removing such cases.

### 11 Apr 2024 v2

@@ 4.2 Box Least-squares Transit Detection
With the detrended light curves, we perform a box least-squares (BLS; Kovács et al., 2002) search to identify planetary transits in the TESS light curves. We sample 100,000 periods linearly spaced between 0.5 and 20 days and 500 transit durations linearly sampled between 0.008 times and 0.08 times of each tested orbital period.
4.3 Vetting Threshold-crossing Events
@@
We identify any signal with a calculated signal-to-pink noise ratio (Hartman & Bakos, 2016) 8 to be a threshold-crossing event (TCE). For each individual TCE, we perform a visual inspection to vet for false positives resulting from both astrophysical events and detrending artifacts. Each candidate event is inspected to determine if the signal that triggered the TCE is transit shaped. Further, we require a minimum of three transit events in order to determine a period.
@@
We search for the presence of secondary eclipses and differences in the depths between even and odd transits to determine if the transit is from an obvious eclipsing binary (EB). For identified TCEs, we fit for the transit depth of even and odd transits respectively to rule out EB scenarios. To measure the transit depth, we fix all parameters except for transit depth, . The transits are modeled using batman (Kreidberg, 2015). We explore the best-fit transit depth and posteriors with a Markov Chain Monte Carlo run implemented with emcee (Foreman-Mackey et al., 2013). We require the difference in our calculated odd and even transit depths to be to not be classified as an EB, though no reported candidate has an odd-even difference in our sample. We visually vet for secondary eclipses by phase folding the detrended TESS light curves at the derived TCE period and transit time. The entire phase-folded light curve is visually inspected for signs of secondary eclipses. Targets exhibiting EB-like secondary eclipses, including those in eccentric orbits, are excluded as planet candidates.
@@
We check for obvious nearby eclipsing binaries (NEBs) by comparing the pixel-by-pixel light curves centered around the target star to determine if the signal that triggered the TCE is consistent with being on target. We then perform a centroid analysis with the TESS FFIs to create difference images using the transit-diffImage code base333https://github.com/stevepur/transit-diffImage following the techniques described in Bryson et al. (2013). For each candidate in each sector, we produce a difference image between the in-transit and out-of-transit pixel stamps. We calculate the centroid of the target star via the TESS pixel response function fit to the difference image. The calculated centroid is compared against the expected centroid of the target star based on a catalog projection. Some offset in the difference image is expected if nearby stars are present, as the centroid naturally shifts away from the dimming star during transit. We flag any centroid offset pixel for visual inspection to determine if the high centroid offset is due to an astrophysical false positive.
@@
Remaining candidates were then checked for planetary multiplicity. We masked the initial signal that triggered the TCE in the nondetrended light curves and then repeated the detrending and transit search (see Sections 4.1 and 4.2). Any signal that pass our S/N criteria is visually vetted in the same method as mentioned above. We additionally check for aliasing of the original signal recovered. All remaining candidates are compared to the list of TOIs and known planets.
@@
Our stellar sample contained 26 TOIs (see Table 2), 22 of which were within our search parameters (0.5 days 20 days and ), as well as two previously previously confirmed transiting planets from K2. Of these, 16 have both FFI light curves and 2 minute cadence light curves (see Table 2), as well as the two K2 planets. Of the 22 known TOIs within our survey parameters, 20 are either planet candidates or confirmed planets, the remaining two have been ruled as either false alarms or false positives by the TESS Follow-up Observing Program Working Groups (TFOP WG; SG1 and SG2).
@@
We recovered all known TOIs in our planet search using the SPOC light curves, including 15 planet candidates within our search parameters and the known planets V1298 Tau c and d. Our QLP planet search recovered all 20 TOIs within our parameters but was unable to recover V1298 Tau c and d. Our pipeline identified a signal corresponding to V1298 Tau c, with a signal-to-pink noise ratio of 6.29, which did not trigger a TCE. Upon visual inspection of the FFI light curves for V1298 Tau, the long cadence of the FFIs and the heightened stellar activity subsequently diluted the transit signal in our detection pipeline.
@@
We identified new planet candidates TIC 88785435.01, TIC 150070085.01, TIC 434398831.01, and TIC 434398831.02 (S. Vach et al. 2024, in preparation) in our QLP search. Out of the new planet candidates, only TIC 150070085.01 also has SPOC data and is independently identified by our SPOC search. These candidates are described in detail in Appendix C.
| TIC | Name | () | (days) | Pipeline | (K) | Age (Myr) | Association | False-positive Rate |
@@
| 299798795 | TOI-1224.01 | 1.900.050 | 4.18 | SPOC and QLP | 3390 | 178 | MELANGE-5 | 0 |
| 299798795 | TOI-1224.02 | 3.000.14 | 17.95 | SPOC and QLP | 3390 | 178 | MELANGE-5 | 0 |
@@
| 384984325 | TOI-6109.01 | 4.510.46 | 8.54 | SPOC and QLP | 5760 | 55 | Theia133 | 0 |
| 384984325 | TOI-6109.02 | 2.780.32 | 5.70 | SPOC and QLP | 5760 | 55 | Theia133 | 0 |
| 410214986 | DS Tuc Ab | 5.700.17 | 8.14 | SPOC and QLP | 5410 | 40 | Tuc-Hor | 0 |
| 441420236 | AU Mic b | 4.200.20 | 8.46 | SPOC and QLP | 3590 | 25 | Pic | 0 |
| 441420236 | AU Mic c | 3.240.16 | 18.86 | SPOC and QLP | 3590 | 25 | Pic | 0 |
| 460205581 | TOI-837 b | 8.601.0 | 8.32 | SPOC and QLP | 6100 | 35 | IC2602 | 0 |
| 464646604 | HIP 94235 b | 3.000.21 | 7.71 | SPOC and QLP | 5990 | 120 | AB Dor | 0 |
@@
| 257605131 | TOI-451 b | 1.86 | SPOC and QLP | 5480 | 120 | Pisces-Eridanus | 0 | |
| 257605131 | TOI-451 c | 9.19 | SPOC and QLP | 5480 | 120 | Pisces-Eridanus | 0 | |
| 257605131 | TOI-451 d | 16.36 | SPOC and QLP | 5480 | 120 | Pisces-Eridanus | 0 | |
@@
| 46631742 | TOI-5358.01 | 2.610.14 | 2.66 | SPOC and QLP | 4640 | 134 | Theia369 | 0.011260.00088 |
| 157081737 | TOI-6095.01 | 1.870.16 | 2.16 | SPOC and QLP | 5260 | 106 | Theia301 | 0.12040.0020 |
@@
| 64837857 | TOI-6550.01 | 14.1 | 20.3 | SPOC and QLP | 6740 | 43 | Theia 117 | |
| 86951294 | TOI-4596.01 | 2.72 | 4.12 | SPOC and QLP | 5600 | 55 | Theia 133 | |
| 166527623 | HIP 67522 b | 10.1 | 6.96 | SPOC and QLP | 5770 | 17 | Sco-Cen | |
@@
| 117689799 | TOI-3504.01 | 14.2 | 2.16 | SPOC and QLP | 5860 | 99 | Theia 254 | |
| 133505138 | TOI-4359.01 | 2.47 | 2.436 | SPOC and QLP | 6530 | 91 | Theia 215 | |
| 164150539 | TOI-1372.01 | 8.22 | 6.16 | SPOC and QLP | 5770 | 100 | Theia 312 | |
| 294500964 | TOI-2550.01 | 10.5 | 9.07 | SPOC and QLP | 5740 | 83 | Theia 211 | |
| 457939414 | TOI-1990.01 | 11.8 | 3.78 | SPOC and QLP | 6080 | 35 | Theia 92 | |
David et al. (2019); P. Thao et al. (2024, in preparation); Wood et al. (2023); A. Dattilo et al. (2024; in preparation); Newton et al. (2019); Plavchan et al. (2020); Bouma et al. (2020); Zhou et al. (2022); This work; S. Vach et al. (2024; in preparation); Newton et al. (2021); () Rizzuto et al. (2020); () Zakhozhay et al. (2022);
@@
4.4 False-positive Rate
Our planet detection pipeline recovered 14 confirmed planets within our SPOC parent sample, and 16 confirmed planets in the QLP parent sample. For planets with confirmed planet dispositions by the TFOP WGs, we adopt a false-positive rate of zero, as these planets have undergone extensive spectroscopic and photometric follow-up observations to rule out false-positive scenarios.
Five remaining planet candidates have not received sufficient follow up to be categorized as verified or validated planets. As such, there exists a possibility that they are the result of astrophysical false-positive scenarios, instrumental systematics, stellar variability, detrending artifacts, etc. There are insufficient planet candidates around young stars from which we can derive a secure false-positive rate estimate.
To attain false-positive estimates, we make use of the publically available TRICERATOPS (Giacalone et al., 2021) code to estimate the probability that a transit signal is planetary in nature or the result of an astrophysical false positive. We explore the impact on the false-positive rates when derived using blend analyses codes, including TRICERATOPS and MOLUSC (Wood et al., 2021). The false-positive rates from TRICERATOPS are adopted in our occurrence rate calculations hereafter and are typically 1-5% for a given candidate. When we adopt TRICERATOPS and MOLUSC together, the false-positive rate approaches 0% for all candidates, and its impact on the calculated occurrence rates is the same as if we assume all candidates are true planets. This upper-bound case is further explored in Appendix B. We do not attempt to differentiate the false-positive rates between multi- and single-planet systems. We present our calculated false-positive rates in Table 2.
We adopt a false-positive rate of 0% for TIC 434398831.01 and .02, which are presented in this work. These are characterized as confirmed planets as space- and ground-based follow up have ruled out any false-positive scenarios (see Appendix C).
@@ 4.4 False-positive Rate
4.5 Candidates Identified but Not Included in the Occurrence Rate Calculations
We identify nine planets or planet candidates that are not included in our occurrence rate calculations. These are presented in Table 2. HIP 67522 b (, Rizzuto et al., 2020), HD 114082 (single transit, , Zakhozhay et al., 2022), and TOI-6550.01 (, days) were recovered in both our SPOC and QLP pipelines, but are not included in our occurrence rate calculations as they fall outside the bounds our survey investigates. We note that TOI-2519.01 (QLP only) and TOI-4596.01 (SPOC and QLP), which we recover in our survey, both show no Li features in ground-based follow-up observations. Upon visual inspection, their respective light curves are also not consistent with youth. We therefore exclude them from our occurrence rate calculations as they are likely not members of their respective clusters. TOI-4596 has an additional single transit event, but we are unable to constrain the period.
TOI-6555.01 is excluded as our vetting processes did not find it to be consistent with being on target. Additionally, TOI-6555.01 was not alerted until Sector 65, which was not made available until after 2023 February.
We identify two new additional candidates, TIC 36332984.01 and TIC 238395674.01, which both pass all of our initial vetting procedures, but we are unable to uniquely determine the orbital periods. TIC 238395674 is likely a multiplanet system, where we observe two transits of TIC 238395674.01 and an additional single transit of TIC 238395674.02. Therefore we do not include these candidates in our occurrence rate calculations.
Additionally, TOI-6715, previously identified as PATHOS 31 (Nardiello et al., 2020), has both QLP and SPOC data. However, as of 2023 February only the QLP data were available, therefore we only include TOI-6715 in our QLP occurrence rate calculations, but not in our SPOC occurrence rate calculations.
4.6 False-positive TESS Objects of Interest Not Included in the Occurrence Rate Calculations
@@
Follow-up observations by the TFOP WGs rule five as nearby planet candidates, meaning the signal triggering the TCE is from a planet transiting a nearby star, or ambiguous planet candidates. TFOP WG SG2 identified the remaining two as a NEB and an EB respectively.
@@ 5 Injection Recovery Test
We perform a set of transit injection and recovery tests to test the completeness of the transit detection pipeline. We perform 1000 iterations of transit injections into each of the nondetrended 7154 QLP and 1927 SPOC light curves. The injected transits are modeled via batman (Kreidberg, 2015), with periods, , and radii, , sampled from log-uniform distributions ranging from 0.5 days to 20 days, and 1 to , respectively, impact parameters, , selected from a uniform distribution, between 0 and 0.8, and a transit time, , uniformly selected from a random phase. The models are resampled to 2, 10, and 30 minute cadences, and multiplied into the nondetrended QLP and/or SPOC light curves for each target. The injected light curve is then run through our detection pipeline as described in Section 4.
An injected planet is recovered if the BLS signal-to-pink noise ratio threshold is met and if the output ephemeris matches that of the injection. As with our planet detection pipeline, we require a TCE to have a signal-to-pink noise ratio . The detected TCE is also required to have its detected transit time and period from the BLS to match the injected time and period with a significance of and (see Coughlin et al. (2014) for definitions of and ).
We present our derived injection and recovery maps in Figure 6. These maps present our full series of injection and recovery tests summed across the entire parent stellar population of QLP and SPOC respectively. Our recoverability of larger planets (5) is consistent across the two surveys, but we see a sharper decline in recoverability in the region between 2 and 4 for QLP compared to SPOC. This is consistent with our expectations due to the difference in stellar populations between the two samples. The on-average larger stellar radii in the QLP sample lead to more difficulties in recovering smaller planets.
We also investigate our susceptibility to false alarms with an inverted light-curve test, as described in Appendix A, finding our manual vetting to be sufficient in removing such cases.